Two machines are used for filling plastic bottle with a net volume of 16.0 ounces. The fill volume can be assume normal with standard deviation of q1= 0.020 and q1= 0.025 ounces. A member of quality member staff suspects that both machines fill to the same net volume, wether or not this volume is 16.0 ounces. A random sample of 10 is taken from the output of each machine. Use a=0.10
1
Expert's answer
2021-10-17T18:05:10-0400

"n_1=10 \\

mean = \bar{x_1} = \frac{1667+16.29+...+15.88+16.99}{10} = 16.339 \\

\sigma_1 = 0.020 \\

n_2 = 10 \\

\bar{x_2} = \frac{16.19+16.49+...+16.39+16.91}{10}=16.14 \\

\sigma_2 = 0.025"

Now, if two machines are used for filling plastic bottles with a net volume of 16.0 ounces , then the net mean difference between the two machines would be zero

"H_0: \mu_1 -\mu_2 = 0 \\

H_1: \mu_1 -\mu_2 ≠ 0"

It is a two-tailed test

α=0.10

Critical value

"Z_c = 1.645"

Reject H0 if |Z| > 1.645

Test-statistic:

"Z = \frac{\bar{x_1} - \bar{x_2}}{\sqrt{\frac{\sigma^2_1}{n_1} + \frac{\sigma_2^2}{n_2}}} \\

Z = \frac{16.339-16.14}{\sqrt{\frac{0.02^2}{10} + \frac{0.025^2}{10}}} \\

Z = -19.66"

Here we noticed that "|z| = 19.656 > Z_c = 1.645"

Hence, null hypothesis is rejected

Conclusion: There is enough evidence to claim that the two machines which are used for filling plastic bottles do not have a net volume of 16.0 ounces​, at the "\alpha = 0.1" significance level